Suzie applies for and obtains a $40,000 low-interest loan for the company from the city council, which has recently passed an initiative encouraging business development related to outdoor activities. The loan is due in three years, and 6% annual interest is due each year on July 31.
Answer #1
assuming the question asked is how much she has to repay at the end of the third year
=40000(1+0.06)^3 =$47640.64
hence interest =$7640.64
Answer #2
40000 * (1 + 0.06 ) ^ 3 = 47640.64 dollars